# CECILIA: Gas-Phase Physical Conditions and Multi-Element Chemistry at Cosmic Noon

### v1

@@ Abstract
Galaxies at Cosmic Noon (z2-3) are characterized by rapid star formation that will lead to significant metal enrichment in the interstellar medium (ISM). While much observational evidence suggests that these galaxies are chemically distinct from those in the local Universe, directly measuring the ISM chemistry in large samples of high-z galaxies is only now possible with the observational capabilities of JWST. In this first key paper of the CECILIA program, we present the direct-method physical conditions and multi-element abundances in twenty galaxies at Cosmic Noon. Using a combination of archival Keck/MOSFIRE and new 30-hr NIRSpec spectroscopy, we measure multiple electron gas densities and the temperature structure from the O+ and S2+ ions. We find that ne[O II] and ne[S II] are comparable but elevated with respect to ne in local star-forming galaxies, and the simultaneous Te[O II] and Te[S III] generally agree with photoionization model Te scaling relations. The O abundances in the CECILIA galaxies range from 12+log(O/H)7.76-8.78 (12-123% solar O/H), representing some of the highest direct-method metallicities and lowest Te measured with JWST to date. The CECILIA galaxies exhibit significantly sub-solar S/O and Ar/O, in agreement with emerging results from other high-z studies and a signature of predominant enrichment from core collapse supernovae. The N/O-O/H trends in the CECILIA galaxies generally agree with the abundance trends in local nebulae, but the large scatter in N/O could be sensitive to the star-formation history. The CECILIA observations underscore the need for exceptionally deep spectroscopy to unveil the ISM abundance patterns in high-z galaxies.
@@ 1 Introduction
The evolving chemistry of galaxies is key to understanding many outstanding questions in contemporary astrophysics. The composition of protostellar gas affects the initial mass function, multiplicity, and planetary populations of stellar systems. The metallicity of the massive stars formed from this gas affects their evolution and ionizing photon production, which in turn impacts the timescale of cosmic reionization. Further, these same massive stars produce transients and gravitational-wave events, and stellar metallicities offer a powerful way to differentiate between progenitor models. Likewise, determining chemical abundance patterns advances our understanding of galaxy star formation histories, gas accretion and outflows, and the circum- and intergalactic media.
@@
For these reasons, reconstructing metallicities and chemical abundance patterns across cosmic time is a high-priority goal for the astronomical community. This is especially the case for galaxies during the epoch of Cosmic Noon, the period at z2-3 when the cosmic star-formation rate (SFR) density peaked (e.g., P. Madau & M. Dickinson, 2014). Given the rapid star formation taking place in these galaxies, it is unsurprising that they show properties distinct from their local counterparts. For example, large ground-based surveys have found that the typical star-forming galaxy (SFG) at 2z3 shows elevated emission from collisionally-excited lines (CELs) of metals relative to local SFGs (e.g., D. Masters et al., 2014; C. C. Steidel et al., 2014; A. E. Shapley et al., 2015). The increased CEL emission has been attributed to ionization from a harder spectrum, likely due to low-metallicity (i.e., low Fe/H) stellar populations at fixed gas-phase O/H (C. C. Steidel et al., 2016; A. L. Strom et al., 2017; A. E. Shapley et al., 2019; L. Clarke et al., 2023). This O/Fe enhancement has been corroborated with observations of the rest-UV spectrum of SFGs at Cosmic Noon, which are best reproduced with low-metallicity stellar continua (C. C. Steidel et al., 2016; F. Cullen et al., 2019; D. Kashino et al., 2022; N. Chartab et al., 2024). Simultaneous observations and photoionization modeling of the rest-UV continuum and rest-optical metal CELs indicates that the relative abundance of O/Fe in these galaxies is enhanced by more than a factor of two relative to the solar abundance ratio (A. L. Strom et al., 2018; F. Cullen et al., 2021; T. M. Stanton et al., 2024). O and Fe are sensitive to different enrichment sources, where O enrichment occurs on short timescales via core-collapse supernovae (CCSNe, 10 Myr) while the bulk of Fe enrichment from Type Ia supernovae occurs on significantly longer timescales (0.1 Gyr, C. Kobayashi & K. Nomoto, 2009). It may be expected, therefore, that the abundance ratios like O/Fe evolve significantly during Cosmic Noon owing to the intense star formation and overall abbreviated star-formation histories (SFHs).
@@
These distinct chemical abundance patterns affect the ionizing spectrum of the massive stars considerably (e.g., J. J. Eldridge & E. R. Stanway, 2009), but they also fundamentally shape the physical conditions in the Interstellar Medium (ISM), specifically the electron gas temperature (Te) and density (ne). Optical emission from forbidden metal CELs is an efficient coolant in the ISM (D. E. Osterbrock & G. J. Ferland, 2006), such that altering the overall metallicity and relative chemical abundance patterns may change the cooling structure within the ISM. Ground-based surveys have also provided evidence of higher ISM ne during Cosmic Noon relative to local SFGs (K. N. Hainline et al., 2009; F. Bian et al., 2010; R. L. Sanders et al., 2016). High ne can significantly decrease CEL emission through collisional de-excitation, producing CEL ratios that mirror those produced in extremely metal-poor environments (e.g., M. W. Topping et al., 2024). Therefore, understanding the chemical evolution of galaxies at Cosmic Noon requires a complementary appreciation of how the ISM Te and ne respond to these unique chemical abundance patterns.
@@
Metal CEL emission is sensitive to Te, ne, gas-phase metallicity, and the ionization state of the gas, thus the intensity of any one CEL is degenerate with multiple parameters. However, the relative intensity ratio of two CELs from the same metal ion eliminates these degeneracies and is sensitive to Te or ne alone. With Te, ne, and the metal CELs, it is then possible to directly measure the metal abundances in the gas (H. L. Dinerstein, 1990; M. Peimbert et al., 2017; R. Maiolino & F. Mannucci, 2019). The direct abundance technique presents an opportunity to assess the chemical enrichment of the ISM, but has one major observational hurdle: the rest-optical metal auroral lines required to measure Te are exceptionally faint, typically 1% the intensity of the strong H I recombination lines. As a result, direct method abundances have been largely inaccessible from ground-based spectroscopy of high-z SFGs (see examples in D. P. Stark et al., 2013; B. L. James et al., 2014; D. A. Berg et al., 2018; R. L. Sanders et al., 2020; L. Clarke et al., 2023; A. Citro et al., 2024).
@@
The prospects for applying the direct abundance method at Cosmic Noon have greatly improved in the years following the launch of JWST. Indeed, JWST/NIRSpec observations have produced a windfall of high-quality rest-optical and NIR spectroscopy of distant galaxies with auroral line detections, allowing Te, ne, and chemical abundance patterns to be measured in galaxies during Cosmic Noon (e.g., N. S. J. Rogers et al., 2024; B. Welch et al., 2024, 2025; T. Morishita et al., 2025; E. Cataldi et al., 2025) and at z6 (e.g., K. Z. Arellano-Córdova et al., 2022; D. Schaerer et al., 2022; M. Curti et al., 2023; I. H. Laseter et al., 2024; M. W. Topping et al., 2024; F. Cullen et al., 2025; C. L. Pollock et al., 2025). Despite these gains afforded by JWST spectroscopy, the evolution of ISM physical conditions and chemical abundance patterns of multiple elements remain uncertain. For example, JWST observations have provided further evidence of ISM density increasing with redshift, both as traced by the low-ionization gas of S+ and O+ (Abdurro’uf et al., 2024) and high-ionization gas of C2+ and N3+ (M. Curti et al., 2025; M. W. Topping et al., 2024, 2025a). However, due to the narrow wavelength separation of the [O II]3727,29 doublet and the low spectral resolution of most JWST/NIRSpec data, simultaneous ne[S II] and ne[O II] are often unavailable in high-z galaxies (c.f., S. Li et al., 2025; B. Welch et al., 2025). While ne[S II] and ne[O II] show good agreement in local nebulae (J. E. Méndez-Delgado et al., 2023), the low-ionization potential of S+ (10.6 eV) means that it exists outside the H II region. A departure of ne[S II] from ne[O II], should one exist, would point to a changing density structure in the ISM relative to local galaxies and would provide insight into the star formation taking place at Cosmic Noon.
@@
Additionally, most direct Te measurements from JWST/NIRSpec observations are made using the [O III]4364 auroral line, a consequence of the high ionization and low metallicities typical of high-z SFGs. In the local Universe, Te measured from different ions in the ISM are strongly correlated and parameterized by Te scaling relations (R. C. Kennicutt et al., 2003; C. Esteban et al., 2009; K. V. Croxall et al., 2016; N. S. J. Rogers et al., 2021). These Te scaling relations allow for the inference of Te in different components of the ISM when a direct temperature is unavailable, but the unique abundance patterns in high-z galaxies could significantly affect the Te structure within the ionized gas and change the functional form of the Te scaling relations. Therefore, Te scaling relations represent a significant systematic uncertainty for chemical abundance surveys that cannot directly resolve the complex Te structure of the multi-phase ISM (K. Z. Arellano-Córdova & M. Rodríguez, 2020; N. S. J. Rogers et al., 2022).
@@
Finally, JWST observations have verified the distinct chemical abundance patterns in high-z SFGs relative to SFGs in the local Universe, both in terms of the bulk metallicity and relative abundance patterns. Direct abundance measurements in galaxies at z2 have revealed unexpected trends such as N-rich, O-poor galaxies at z2 (B. Welch et al., 2024, 2025; K. Z. Arellano-Córdova et al., 2025) and super-solar N/O at z6 (P. Senchyna et al., 2024; M. W. Topping et al., 2024, 2025b; M. Curti et al., 2025). Additionally, there is growing evidence to suggest that Ar/O is sub-solar in high-z galaxies (N. S. J. Rogers et al., 2024; S. Bhattacharya et al., 2025a; T. M. Stanton et al., 2025), which is unexpected if both elements originate primarily from CCSNe. The abundance patterns of other elements, such as S, remain largely unconstrained at high-z (c.f., B. Welch et al., 2025; T. Morishita et al., 2025) owing to small sample sizes and the faint emission lines required for the direct abundance method. Therefore, it remains unclear if these peculiar abundance patterns are representative of the typical SFG at Cosmic Noon, or if they are a reflection of prompt enrichment from recent phases of intense star formation (P. Senchyna et al., 2024; C. Kobayashi & A. Ferrara, 2024).
@@
Uncovering the evolution of the ISM physical conditions, chemical abundance patterns, and the prevalence of different enrichment mechanisms, therefore, requires a large sample of high-z SFGs with direct-method metallicities. In this paper, we present such a sample, acquired as part of the Chemical Evolution Constrained using Ionized Lines in Interstellar Aurorae (CECILIA, A. L. Strom et al., 2023) program. CECILIA was designed to detect the Te-sensitive auroral lines of [O II] and [S III] in SFGs at Cosmic Noon using ultra-deep NIRSpec observations and complementary Keck/MOSFIRE spectroscopy of the strong optical CELs. These deep data have enabled studies of the faintest emission lines observed in the rest-optical and NIR spectra of high-z galaxies (A. L. Strom et al., 2023), the chemical abundance patterns in the ISM at z3 (N. S. J. Rogers et al., 2024), and the ionization conditions and star formation in low-mass, faint galaxies at Cosmic Noon (M. Raptis et al., 2025). We now examine the full sample of ne, Te, and chemical abundances measured in CECILIA, representing a significant increase in the number of galaxies with direct abundances at Cosmic Noon.
@@
This manuscript is organized as follows: we describe the CECILIA survey, MOSFIRE and NIRSpec data reduction, emission line fitting, and reddening corrections in §2. In §3, we outline the methods used for ne and Te calculations in the CECILIA galaxies, and we discuss the simultaneous ne[O II]-ne[S II] and Te[O II]-Te[S III] trends measured at Cosmic Noon. In §4, the chemical abundance trends in the full sample are presented and analyzed. We summarize our conclusions in §5, and we elaborate on specific aspects of the data reduction and analysis in Appendices A and B. Throughout this paper, we assume a CDM cosmology with = 70 km/s/Mpc, = 0.7, and = 0.3. Solar metallicities are adopted from M. Asplund et al. (2021): 12+log(O/H)⊙ 8.690.04, log(N/O)⊙ 0.860.08, log(S/O)⊙ 1.570.05, and log(Ar/O)⊙ 2.310.11. We refer to the position of spectral features using their vacuum wavelengths. When discussing individual galaxies in the text, we use the galaxy ID and omit the ”Q2343-” field name.
@@ 2.1 Galaxy Sample
CECILIA includes galaxies that were first identified and observed as part of the Keck Baryonic Structure Survey (KBSS; C. C. Steidel et al. 2010; G. C. Rudie et al. 2012; A. L. Strom et al. 2017). KBSS is a large spectroscopic survey of galaxies in 15 fields surrounding bright quasars; most KBSS galaxies are selected based on their rest-UV continuum colors, with a subset of continuum-faint galaxies selected on the basis of their Ly emission in narrow-band imaging. The new JWST/NIRSpec observations were conducted in the Q2343+125 field due to the (1) high density of sources with spectroscopic redshifts, (2) large catalog of narrow-band-selected Ly emitters (LAEs) at , and (3) existing HST/WFC3 F140W mosaic that provided precision astrometry for mask design and galaxy size measurements. The ancillary data available for the Q2343+125 field and, by extension, for CECILIA include , , and NIR (rest-optical) spectroscopy from Keck/MOSFIRE, deep optical (rest-UV) spectroscopy from Keck/LRIS, and imaging in through , F140W, F160W, IRAC Ch , MIPS 24 m, and narrow-band (rest-frame) Ly filters.
Given the explicit goal of measuring electron temperatures and direct-method metallicities from the JWST/NIRSpec observations, we prioritized galaxies with , where all of the required emission lines fall within the G235M, G395M, and ground-based spectral bandpasses. As described in more detail in A. L. Strom et al. (2023), galaxies were also assigned a higher priority if they have smaller than average sizes, high star formation rates (SFRs), and/or large observed nebular [O III] line fluxes, all of which increase the ease of detecting the faint temperature-sensitive auroral lines. The full targeted sample spans the parameter space occupied by typical SFGs in various line ratio diagrams such as the N2-BPT (J. A. Baldwin et al., 1981) and - (or ionization-excitation diagram) in order to provide a useful metric of the ISM properties across the parent galaxy population. We also included a number of narrow-band selected Ly emitters (LAEs) from R. F. Trainor et al. (2015) with spectroscopic detections of Ly and [O III] or H; the first results for these faint galaxies are presented in M. Raptis et al. (2025) and are not the main focus of the present work.

### v2

@@ Abstract
Galaxies at Cosmic Noon (z2-3) are characterized by rapid star formation that will lead to significant metal enrichment in the interstellar medium (ISM). While much observational evidence suggests that these galaxies are chemically distinct from those in the local Universe, directly measuring the ISM chemistry in large samples of high-z galaxies is only now possible with the observational capabilities of JWST. In this first key paper of the CECILIA program, we present the direct-method physical conditions and multi-element abundances in twenty galaxies at Cosmic Noon. Using a combination of archival Keck/MOSFIRE and new 30-hr NIRSpec spectroscopy, we measure multiple electron gas densities and the temperature structure from the O+ and S2+ ions. We find that ne[O II] and ne[S II] are comparable but elevated with respect to ne in local star-forming galaxies, and the simultaneous Te[O II] and Te[S III] generally agree with photoionization model Te scaling relations. The O abundances in the CECILIA galaxies range from 12+log(O/H)7.76-8.81 (12-131% solar O/H), representing some of the highest direct-method metallicities and lowest Te (Te[O II]6500 K) measured with JWST to date. The CECILIA galaxies exhibit significantly sub-solar S/O and Ar/O a signature of predominant enrichment from core collapse supernovae. The N/O-O/H trends in the CECILIA galaxies generally agree with the abundance trends in local nebulae, but the large scatter in N/O could be sensitive to the star-formation history. The CECILIA observations demonstrate that exceptionally deep JWST spectroscopy can unveil the multi-element ISM abundance patterns in typical high-z galaxies.
@@ 1 Introduction
The evolving chemistry of galaxies is key to understanding many outstanding questions in contemporary astrophysics. The composition of protostellar gas affects the initial mass function, multiplicity, and planetary populations of stellar systems (e.g., badenes2018; moe2019; sharda2022). The metallicity of the massive stars formed from this gas affects their evolution and ionizing photon production, which in turn impacts the timescale of cosmic reionization (e.g., stanway2016; gotberg2017) and the rates of transient and gravitational-wave events (e.g., eldridge2019; neijssel2019; vanson2025). Likewise, determining chemical abundance patterns advances our understanding of galaxy star formation histories, gas accretion and outflows, and the circum- and intergalactic media (e.g., tinsley1979; mcwilliam1997).
@@
For these reasons, reconstructing metallicities and chemical abundance patterns across cosmic time is a high-priority goal for the astronomical community. This is especially the case for galaxies during the epoch of Cosmic Noon, the period at z2-3 when the cosmic star-formation rate (SFR) density peaked (e.g., madau2014). Given the rapid star formation taking place in these galaxies, it is unsurprising that they show properties distinct from their local counterparts. For example, large rest-optical surveys using ground-based NIR spectrographs have found that the typical star-forming galaxy (SFG) at 2z3 shows elevated emission from collisionally-excited lines (CELs) of metals relative to local SFGs (e.g., masters2014; steidel2014; shapley2015). The increased CEL emission has been attributed to ionization from a harder spectrum, likely due to low-metallicity (i.e., low Fe/H) stellar populations at fixed gas-phase O/H (steidel2016; strom2017; shapley2019; clarke2023). This O/Fe enhancement has been corroborated with observations of the rest-UV spectrum of SFGs at Cosmic Noon, which are best reproduced with low-metallicity stellar continua (steidel2016; cullen2019; kashino2022; chartab2024). Simultaneous observations and photoionization modeling of the rest-UV continuum and rest-optical metal CELs indicates that the relative abundance of O/Fe in these galaxies is enhanced by more than a factor of two relative to the solar abundance ratio (strom2018; topping2020a; topping2020b; cullen2021; stanton2024). O and Fe are sensitive to different enrichment sources, where O enrichment occurs on short timescales via core-collapse supernovae (CCSNe, 10 Myr) while the bulk of Fe enrichment from Type Ia supernovae occurs on significantly longer timescales (0.1 Gyr, kobayashi2009). It may be expected, therefore, that abundance ratios like O/Fe evolve significantly during Cosmic Noon owing to the intense star formation and overall abbreviated star-formation histories (SFHs).
@@
These distinct chemical abundance patterns affect the ionizing spectrum of the massive stars considerably (e.g., eldridge2009), but they also fundamentally shape the physical conditions in the Interstellar Medium (ISM), specifically the electron gas temperature (Te) and density (ne). Optical emission from forbidden metal CELs is an efficient coolant in the ISM (osterbrock2006), such that altering the overall metallicity and relative chemical abundance patterns may change the cooling structure within the ISM. Ground-based surveys have also provided evidence of higher ISM ne during Cosmic Noon relative to local SFGs (hainline2009; bian2010; sanders2016). High ne can significantly decrease CEL emission through collisional de-excitation, producing CEL ratios that mirror those produced in extremely metal-poor environments (e.g., topping2024). Therefore, understanding the chemical evolution of galaxies at Cosmic Noon requires a complementary appreciation of how the ISM Te and ne respond to these unique chemical abundance patterns.
@@
Metal CEL emission is sensitive to Te, ne, gas-phase metallicity, and the ionization state of the gas, thus the intensity of any one CEL is degenerate with multiple parameters. However, the relative intensity ratio of two CELs from the same metal ion eliminates these degeneracies and is sensitive to Te or ne alone. With Te, ne, and the metal CELs, it is then possible to directly measure the metal abundances in the gas (dinerstein1990; peimbert2017; maiolino2019). The direct abundance technique presents an opportunity to assess the chemical enrichment of the ISM, but has one major observational hurdle: the rest-optical metal auroral lines required to measure Te are exceptionally faint, typically 1% the intensity of the strong H I recombination lines. As a result, direct method abundances have been largely inaccessible from ground-based spectroscopy of high-z SFGs (see examples in stark2013; james2014; berg2018; sanders2020; clarke2023; citro2024).
@@
The prospects for applying the direct abundance method at Cosmic Noon have greatly improved in the years following the launch of JWST. Indeed, JWST/NIRSpec observations have produced a windfall of high-quality rest-optical and NIR spectroscopy of distant galaxies with auroral line detections, allowing Te, ne, and chemical abundance patterns to be measured in galaxies during Cosmic Noon (e.g., rogers2024; welch2024; welch2025; morishita2025; cataldi2025; scholte2025) and at z6 (e.g., arellano-cordova2022; schaerer2022; curti2023; laseter2024; topping2024; chakraborty2025; cullen2025; pollock2025). Despite these gains afforded by JWST spectroscopy, the evolution of ISM physical conditions and chemical abundance patterns of multiple elements remain uncertain. For example, JWST observations have provided further evidence of ISM density increasing with redshift, both as traced by the low-ionization gas of S+ and O+ (abdurrouf2024) and high-ionization gas of C2+ and N3+ (curti2025; topping2024; topping2025). However, due to the narrow wavelength separation of the [O II]3727,29 doublet and the low spectral resolution of most JWST/NIRSpec data, simultaneous ne[S II] and ne[O II] are often unavailable in high-z galaxies (c.f., li2025; welch2025). While ne[S II] and ne[O II] show good agreement in local nebulae (mendez-delgado2023), the low-ionization potential (IP) of S+ (10.6 eV) means that it exists outside the H II region. A departure of ne[S II] from ne[O II], should one exist, would point to a changing density structure in the ISM relative to local galaxies and would provide insight into the star formation taking place at Cosmic Noon.
@@
Additionally, most direct Te measurements from JWST/NIRSpec observations are made using the [O III]4364 auroral line, a consequence of the high ionization and low metallicities typical of high-z SFGs. In the local Universe, Te measured from different ions in the ISM are strongly correlated and parameterized by Te scaling relations (kennicutt2003; esteban2009; croxall2016; rogers2021). These Te scaling relations, either calibrated with empirical Te data or photoionization model predictions (campbell1986; garnett1992; pagel1992), allow for the inference of Te in different components of the ISM when a direct temperature is unavailable. However, the unique abundance patterns in high-z galaxies could significantly affect the Te structure within the ionized gas and change the functional form of the Te scaling relations. Therefore, Te scaling relations represent a significant systematic uncertainty for chemical abundance surveys that cannot directly resolve the complex Te structure of the multi-phase ISM (arellano-cordova2020; rogers2022).
@@
Finally, JWST observations have verified the distinct chemical abundance patterns in high-z SFGs relative to SFGs in the local Universe, both in terms of the bulk metallicity and relative abundance patterns. Direct abundance measurements in galaxies at z2 have revealed unexpected trends such as N-rich, O-poor galaxies at z2 (welch2024; welch2025; arellano-cordova2025) and super-solar N/O at z6 (senchyna2024; topping2024; topping2025no; curti2025). Additionally, there is growing evidence to suggest that Ar/O is sub-solar in high-z galaxies (rogers2024; bhattacharya2025A; stanton2025). Since Ar is produced in Type Ia SNe (kobayashi2020SNIa), the Ar/O ratio should be sensitive to the relative enrichment from CC and Type Ia SNe and, therefore, the SFH of a galaxy. Such a dependence may produce different Ar/O abundance patterns in high-z galaxies compared to typical z0 SFGs with solar Ar/O (e.g., berg2020; arellano-cordova2024a). The abundance patterns of other elements, such as S, remain largely unconstrained at high-z (c.f., welch2025; morishita2025) owing to small sample sizes and the faint emission lines required for the direct abundance method. Therefore, it remains unclear if these peculiar abundance patterns are representative of the typical SFG at Cosmic Noon, or if they are a reflection of prompt enrichment from recent phases of intense star formation (senchyna2024; kobayashi2024).
@@
Uncovering the evolution of the ISM physical conditions, chemical abundance patterns, and the prevalence of different enrichment mechanisms, therefore, requires a large sample of high-z SFGs with direct-method metallicities. In this paper, we present such a sample, acquired as part of the Chemical Evolution Constrained using Ionized Lines in Interstellar Aurorae (CECILIA, PID2593, strom2023) program. CECILIA was designed to detect the Te-sensitive auroral lines of [O II] and [S III] in SFGs at Cosmic Noon using ultra-deep NIRSpec observations and complementary Keck/MOSFIRE spectroscopy of the strong optical CELs. These deep data have enabled studies of the faintest emission lines observed in the rest-optical and NIR spectra of high-z galaxies (strom2023), the chemical abundance patterns in the ISM at z3 (rogers2024), and the ionization conditions and star formation in low-mass, faint galaxies at Cosmic Noon (raptis2025). We now examine the full sample of ne, Te, and chemical abundances measured in CECILIA, including 9 galaxies with direct Te[S III] and 17 with Te[O II], representing a 70% and 44% increase in the number of galaxies with direct Te[S III] and Te[O II], respectively, at z1.
@@
This manuscript is organized as follows: we describe the CECILIA survey, MOSFIRE and NIRSpec data reduction, emission line fitting, and reddening corrections in §2. In §3, we outline the methods used for ne and Te calculations in the CECILIA galaxies, and we discuss the simultaneous ne[O II]-ne[S II] and Te[O II]-Te[S III] trends measured at Cosmic Noon. In §4, the chemical abundance trends in the full sample are presented and analyzed. We summarize our conclusions in §5, and we elaborate on specific aspects of the data reduction and analysis in Appendices A and B. Throughout this paper, we assume a CDM cosmology with = 70 km/s/Mpc, = 0.7, and = 0.3. Solar metallicities are adopted from asplund2021: 12+log(O/H)⊙ 8.690.04, log(N/O)⊙ 0.860.08, log(S/O)⊙ 1.570.05, and log(Ar/O)⊙ 2.310.11. We refer to the position of spectral features using their vacuum wavelengths. When discussing individual galaxies in the text, we use the galaxy ID and omit the ”Q2343-” field name.
@@ 2.1 Galaxy Sample
CECILIA includes galaxies that were first identified and observed as part of the Keck Baryonic Structure Survey (KBSS; steidel2010; rudie2012; strom2017). KBSS is a large spectroscopic survey of galaxies in 15 fields surrounding bright quasars; most KBSS galaxies are selected based on their rest-UV continuum colors, with a subset of continuum-faint galaxies selected on the basis of their Ly emission in narrow-band imaging. The new JWST/NIRSpec observations were conducted in the Q2343+125 field due to the (1) high density of sources with spectroscopic redshifts, (2) large catalog of narrow-band-selected Ly emitters (LAEs) at , and (3) existing HST/WFC3 F140W mosaic that provided precision astrometry for mask design and galaxy size measurements. The ancillary data available for the Q2343+125 field and, by extension, for CECILIA include , , and NIR (rest-optical) spectroscopy from Keck/MOSFIRE, deep optical (rest-UV) spectroscopy from Keck/LRIS, and imaging in through , F140W, F160W, IRAC Ch , MIPS 24 m, and narrow-band (rest-frame) Ly filters.
Given the explicit goal of measuring electron temperatures and direct-method metallicities from the JWST/NIRSpec observations, we prioritized galaxies with , where all of the required emission lines fall within the G235M, G395M, and ground-based spectral bandpasses. As described in more detail in strom2023, galaxies were also assigned a higher priority if they have smaller than average sizes, high SFRs, and/or large observed nebular [O III] line fluxes, all of which increase the ease of detecting the faint temperature-sensitive auroral lines. The full targeted sample spans the parameter space occupied by typical SFGs in various line ratio diagrams such as the N2-BPT (baldwin1981) and - (or ionization-excitation diagram) in order to provide a useful metric of the ISM properties across the parent galaxy population. We also included a number of narrow-band selected Ly emitters (LAEs) from trainor2015 with spectroscopic detections of Ly and [O III] or H; the first results for these faint galaxies are presented in raptis2025; raptis2025b and are not the main focus of the present work.